# Pressure, Reference Points, and Risk-Taking Behavior: Evidence from Bench Press Competitions

### v1

@@ 7.2 Expected rivals’ attempt weights
In our theoretical prediction in Section 3 and empirical exercises, we treat the attempt weight of lifter $i$’s higher rival, which is not observable to lifer $i$ in his attempt stage, as the actual attempt weight in a rational expectation manner. Although this operation is ad hoc, we believe that treating the higher rival’s attempt weight as completely unknown information ignores pressure from the higher rank rival, which matters in the actual competition.
Table 10 presents the Root Mean Squared Error (RMSE) and $R^{2}$ values for various prediction models of attempt weights using linear regression. Predictive accuracy is assessed through 5-fold cross-validation. All models include gender, body weight, number of competition experiences, a dummy variable for first participation, current best, and fixed effects for equipment category, age class, division, weight class, and federation as predictors. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt—which is unobservable when predicting the higher rival’s 2nd attempt weight—is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of both the 1st and 2nd attempts are used, whereas in the “without 2nd success” model, the success of the 2nd attempt—unobservable when predicting the higher rival’s 3rd attempt weight—is excluded.
The prediction for the 1st attempt yields an RMSE of 24.794 and an $R^{2}$ of 0.798, indicating that accurately predicting the 1st attempt is relatively challenging. In contrast, predictions for the 2nd and 3rd attempts across all models show an $R^{2}$ exceeding 0.99, demonstrating that high predictive accuracy can be achieved even without incorporating the outcome of the preceding attempt, which is unobservable when predicting the attempt weight of a higher-ranked rival. When comparing the full models with the “without” models for both the 2nd and 3rd attempts, the full model exhibits slightly better predictive accuracy. This suggests that the outcome of an attempt reflects the lifter’s condition on the day of the competition, and lifters choose their weights accordingly, while lower-ranked rivals are unable to observe this condition. These results support our assumption that the pressure variable, derived from predictions of the higher rival’s attempt weight, is independent of unobserved factors such as the lifter’s condition and can be considered exogenous.
@@
1st Full 24.795 0.798
2nd Full 4.257 0.994
2nd Without 1st success 4.858 0.993
3rd Full 3.433 0.996
3rd Without 2nd success 4.158 0.995
@@
Notes: All models include gender, body weight, equipment category, age class, division, weight class, federation, first participation, number of competition experiences, and current best as predictors. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of both the 1st and 2nd attempts are used, while in the “without 2nd success” model, the success of the 2nd attempt is excluded. Predictive accuracy is assessed using 5-fold cross-validation.
@@ 8 Conclusion
Using the data from bench press competitions, our study highlights the critical impact of competitive pressure on weight attempts and success probability which are analogous to lifters’ decision-making and performance. Pressure from rivals, particularly during second attempts, leads to more aggressive weight selections, driven by loss aversion and the desire to avoid a lower rank. In contrast, third attempts are influenced by fatigue and the diminishing potential for rank improvement, resulting in more conservative behavior under pressure from higher-ranked rivals. Counterfactual simulations of no-pressure scenarios reveal that lifters tend to be more conservative in their second attempts but more ambitious in their third, though success probabilities decrease without the motivating influence of competition. These results suggest that while pressure can drive lifters to perform at their best, particularly in earlier attempts, maintaining a focus on personal goals and committing to original plans may be essential for maximizing success in later stages of competition.
As a future direction, structural estimation regarding individual lifters across attempts would be interesting. Each player allocates his stamina to three attempts strategically and dynamically, which is a more prevalent feature in the popular SBD (Squat-Benchpress-Deadlift) competition. Incorporating the sequential structure across the first, second, and third attempts for all players, structural estimation of choosing weight attempt, lifting the weight as a lottery choice, and realizing the gain under their stamina constraint would be an interesting avenue.
Appendix A Additional figures and tables
Data patterns of the weight attempt and success probability other than Figures 2 and 3 in HTML in authors’ Github. Overall, we find that lifters appear to struggle more in the third attempt, especially with larger deviations from their best, indicating a significant drop in success probability compared to the second attempt.

### v2

@@ 7.2 Expected rivals’ attempt weights
In our theoretical prediction in Section 3 and empirical exercises, we use the predicted attempt weight of lifter $i$’s higher-ranked rival—unobservable to lifter $i$ at the attempt stage. Although this approach involves prediction errors, assuming that the higher rival’s attempt weight is completely unknown would ignore the pressure exerted by the higher-ranked rival, which plays a significant role in actual competition.
@@
1st Full 23.837 0.813
2nd Full 4.214 0.994
2nd Without 1st success 4.481 0.994
3rd Full 3.450 0.996
3rd Without 2nd success 4.191 0.995
@@
Notes: All models include gender, body weight, equipment category, age class, division, weight class, federation, a dummy for first participation, number of competition experiences, and current best as predictors. They also include a dummy variable indicating whether the lifter shares the same declared attempt weight with another competitor—potentially causing ambiguity in the lifting order due to unresolved tie-breaking rules—and another dummy variable capturing mismatches between the actual lifting order and the interim ranking order, which may affect the set of observable information about rivals. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of the 2nd attempts are used, while in the “without 2nd success” model, the success of the 2nd attempt is excluded. Predictive accuracy is assessed using 5-fold cross-validation.
Table 11 presents the Root Mean Squared Error (RMSE) and $R^{2}$ values for different prediction models of attempt weights using linear regression. Predictive accuracy is assessed through 5-fold cross-validation. All models include gender, body weight, equipment category, age class, division, weight class, federation, a dummy for first participation, number of competition experiences, and current best as predictors. They also include a dummy variable indicating whether the lifter shares the same declared attempt weight with another competitor—potentially causing ambiguity in the lifting order due to unresolved tie-breaking rules—and another dummy variable capturing mismatches between the actual lifting order and the interim ranking order, which may affect the set of observable information about rivals. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt—unobservable when predicting the higher rival’s 2nd attempt weight—is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of both the 1st and 2nd attempts are used, whereas in the “without 2nd success” model, the success of the 2nd attempt—unobservable when predicting the higher rival’s 3rd attempt weight—is excluded.
The prediction for the 1st attempt yields an RMSE of 23.837 and an $R^{2}$ of 0.813, indicating that accurately predicting the 1st attempt remains challenging. In contrast, the predictions for the 2nd and 3rd attempts achieve an $R^{2}$ exceeding 0.99 across all models, demonstrating that even without incorporating the outcome of the preceding attempt, which is unobservable when predicting the higher rival’s attempt weight, high predictive accuracy can still be maintained. Comparing the full models with the “without” models for both the 2nd and 3rd attempts, the full models exhibit slightly lower RMSEs (4.214 vs. 4.481 for the 2nd attempt; 3.450 vs. 4.191 for the 3rd attempt), indicating marginally better predictive accuracy. This suggests that the outcome of an attempt reflects the lifter’s condition on the day of the competition, influencing weight selection, while lower-ranked rivals are unable to observe this condition. These findings support our assumption that the pressure variable—derived from predictions of the higher rival’s attempt weight—is independent of unobserved factors such as the lifter’s condition and can be considered exogenous.
Table 12: Regression of the Difference between the Second and Third Attempt Weight Compared to the Current Best on Pressure with Bootstrap
(1) (2)
Dependent Variable $W_{it}^{2}-$(best) $W_{it}^{3}-$(best)
Male -1.477 3.210
(0.384) (0.529)
Body weight 0.274 0.355
(0.014) (0.033)
Num experience 0.229 0.255
(0.052) (0.098)
1(first participation) 32.849 34.061
(0.989) (0.984)
Pressure, lower rival, 2nd 0.111
Pressure, higher rival, 2nd 0.440
Pressure, lower rival, 3rd 0.068
Pressure, higher rival, 3rd 0.407
Control X X
Num.Obs. 246521 246521
R2 0.536 0.356
R2 Adj. 0.534 0.353
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from a lower rival at the 2nd attempt is calculated as (lower rival 2nd attempt) - (his 1st outcome), and pressure from a lower rival at the 3rd attempt is (lower rival 3rd attempt) - (his 2nd outcome). For higher rivals, predicted attempt weights are obtained as the mean from the bootstrap procedure, while standard errors are computed as the standard deviation of the bootstrap estimates, ensuring that the uncertainty in the prediction process is fully reflected in the inference.
To address the prediction error problem of using predicted values as explanatory variables, we implement a bootstrap procedure when incorporating the predicted attempt weight of lifter $i$’s higher-ranked rival—unobservable to lifter $i$ at the $k$-th attempt stage—into our regression framework. Since this predicted value, denoted by $\hat{Z}_{it}^{k}$, is derived from a separate estimation and contains prediction error, standard errors tend to be downward biased due to the smoothing effect of prediction. To correct for this, we apply a bootstrap approach that resamples the original data at the federation level, re-estimates the prediction model for $\hat{Z}_{it}^{k}$ in each bootstrap iteration, and then regresses the lifter’s own attempt weight or success probability on the re-estimated $\hat{Z}_{it}^{k}$. This procedure ensures that the uncertainty in $\hat{Z}_{it}^{k}$ is incorporated into the standard error estimates by reflecting its sampling variability. Consequently, the bootstrap-corrected standard errors provide more robust inference on the effect of competitive pressure from higher-ranked rivals on weight selection and performance outcomes.
Compared to Table 5, which uses predicted higher rivals’ attempt weights, Table 12 incorporates a bootstrap procedure to account for uncertainty in predicted higher rival attempt weights. This leads to slightly larger standard errors but consistent point estimates and signs. Notably, the male coefficient in the second attempt shifts from –1.279 to –1.477, and pressure from higher-ranked rivals in the second attempt is stable, reinforcing the robustness of competitive pressure effects. Similarly, the effect of pressure from higher-ranked rivals in the third attempt remains stable (0.407 vs. 0.409 in Table 5). The stability of estimates confirms the validity of the benchmark results even when accounting for prediction uncertainty.
@@ 8 Conclusion
Using data from bench press competitions, our study highlights the impact of competitive pressure on lifters’ attempt selections and success probabilities. Pressure from both lower- and higher-ranked rivals in the second attempt encourages more aggressive weight selections. In the third attempt, lifters continue to take risks under pressure from both directions, with higher-ranked rivals in particular inducing significant upward adjustments in attempt weights.
Heterogeneity analysis reveals that male lifters, experienced competitors, and those with repeated rivalry encounters respond more strongly to pressure. Male lifters take greater risks, while female and less experienced lifters remain more conservative. Familiarity with rivals amplifies responsiveness to competition. While success probabilities tend to decline under upward pressure for all lifters, the effect is notably stronger among women, underscoring the psychological difficulty of overtaking stronger competitors.
Counterfactual simulations show that without competitive pressure, lifters tend to adopt more conservative second attempts and maintain or slightly increase third attempt weights. However, the effects on performance are highly heterogeneous: while some lifters benefit from the absence of pressure, many experience lower success rates and reduced expected outcomes, particularly in the second attempt. These results highlight that external competition plays a dual role—motivating some lifters to perform better while destabilizing others—underscoring the importance of individual sensitivity in shaping performance under pressure.
Future research could explore structural estimation of decision-making across attempts. Lifters strategically allocate their stamina, a key aspect in SBD (Squat-Benchpress-Deadlift) competitions. Modeling this sequential structure—where lifters choose attempt weights, treat each lift as a lottery, and adjust under stamina constraints—could offer deeper insights into risk-taking and performance optimization.
Appendix A Online Appendix (Not for publication)
A.1 Comprehensive Attempt Stage Model
In this section, we present an attempt stage model that simultaneously considers the influence of both the lower and higher rivals on the lifter’s weight declaration. By accounting for both sources of competitive pressure simultaneously, this model provides a more complete representation of how lifters adjust their declared weights in response to the surrounding competition. However, the added complexity makes the model less intuitive to interpret, as the interactions between the two rivals’ effects become more intricate.
We maintain the notation used in Section 3, let $w_{L}$ denote the lower rival’s actual attempt weight which is observable to lifter $i$ during the attempt stage. Also, let $w_{H}$ denote the higher rival’s expected attempt weight under lifter $i$’s rational expectations. Although this is not observable to lifter $i$’s attempt stage, we treat this as highly predictable information to lifter $i$. Additionally, let $y_{L}$, $y_{H}$, and $y_{i}$ represent the best outcome weights from earlier attempts within the competition for the lower rival, the higher rival, and lifter $i$, respectively, which are known in lifter $i$’s current attempt stage. The (strict) inequality conditions $y_{L}<y_{i}<y_{H}$, determined by rank order, and $y_{L}<w_{L}$, $y_{i}<w_{i}$, and $y_{H}<w_{H}$, meaning that the outcome weight must not exceed the corresponding attempt weight, must hold. Although we can assume that $w_{L}<w_{H}$, meaning that the lower rival does not attempt to overtake two higher-ranked rivals as inferred from our data, must hold, we keep generality and allow $w_{L}>w_{H}$. These conditions determine the joint probability of the outcome tuple as shown in Table 13.
Table 13: Joint probability of outcome tuple
Pressure $i$ $L$ $H$ Outcome $i$’s rank Probability
$w_{i}<\min\{w_{L},w_{H}\}$ S S S $w_{i}<\min\{w_{L},w_{H}\}$- $q(w)q(w_{L})q(w_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S S F $w_{i}<\min\{w_{L},y_{H}\}$- $q(w)q(w_{L})(1-q(w_{H}))P(w_{i}<y_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S S F $y_{H}<w_{i}<w_{L}$ K $q(w)q(w_{L})(1-q(w_{H}))P(w_{i}>y_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S F S $y_{L}(<y_{i})<w_{i}<w_{H}$ K $q(w)(1-q(w_{L}))q(w_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S F F $y_{L}(<y_{i})<w_{i}<y_{H}$ K $q(w)(1-q(w_{L}))(1-q(w_{H}))P(w_{i}<y_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S F F $y_{L}<y_{H}<w_{i}$ + $q(w)(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ F S Any $y_{i}<\min\{w_{L},y_{H}\}$- $(1-q(w))q(w_{L})$
$w_{i}<\min\{w_{L},w_{H}\}$ F F Any $y_{L}<y_{i}<y_{H}$ K $(1-q(w))(1-q(w_{L}))$
$w_{L}<w_{i}<w_{H}$ S S S $w_{L}<w_{i}<w_{H}$ K $q(w)q(w_{L})q(w_{H})$
$w_{L}<w_{i}<w_{H}$ S S F $w_{L}<w_{i}<y_{H}(<w_{H})$ K $q(w)q(w_{L})(1-q(w_{H}))P(w_{i}<y_{H})$
$w_{L}<w_{i}<w_{H}$ S S F $\max\{w_{L},y_{H}\}<w_{i}$ + $q(w)q(w_{L})(1-q(w_{H}))P(w_{i}>y_{H})$
$w_{L}<w_{i}<w_{H}$ S F S $y_{L}(<w_{L})<w_{i}<w_{H}$ K $q(w)(1-q(w_{L}))q(w_{H})$
$w_{L}<w_{i}<w_{H}$ S F F $y_{L}<w_{i}<y_{H}(<w_{H})$ K $q(w)(1-q(w_{L}))(1-q(w_{H}))P(w_{i}<y_{H})$
$w_{L}<w_{i}<w_{H}$ S F F $y_{L}<y_{H}<w_{i}$ + $q(w)(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})$
$w_{L}<w_{i}<w_{H}$ F S S $y_{i}<w_{L}<w_{H}$- $(1-q(w))q(w_{L})q(w_{H})P(w_{L}>y_{i})$
$w_{L}<w_{i}<w_{H}$ F S S $w_{L}<y_{i}<w_{H}$ K $(1-q(w))q(w_{L})q(w_{H})P(w_{L}<y_{i})$
$w_{L}<w_{i}<w_{H}$ F S F $w_{L}<y_{i}<y_{H}$ K $(1-q(w))q(w_{L})(1-q(w_{H}))P(w_{L}<y_{i})$
$w_{L}<w_{i}<w_{H}$ F S F $y_{i}<\min\{w_{L},y_{H}\}$- $(1-q(w))q(w_{L})(1-q(w_{H}))P(w_{L}>y_{i})$
$w_{L}<w_{i}<w_{H}$ F F S $y_{L}<y_{i}<w_{H}$ K $(1-q(w))(1-q(w_{L}))q(w_{H})$
$w_{L}<w_{i}<w_{H}$ F F F $y_{L}<y_{i}<y_{H}$ K $(1-q(w))(1-q(w_{L}))(1-q(w_{H}))$
$\max\{w_{L},w_{H}\}<w_{i}$ S Any Any $\max\{w_{L},w_{H}\}<w_{i}$ + $q(w)$
$\max\{w_{L},w_{H}\}<w_{i}$ F S Any $w_{L}<y_{i}<\min\{y_{H},w_{H}\}$ K $(1-q(w))q(w_{L})P(w_{L}<y_{i})$
$\max\{w_{L},w_{H}\}<w_{i}$ F S Any $y_{i}<\min\{w_{L},y_{H}\}$- $(1-q(w))q(w_{L})P(w_{L}>y_{i})$
$\max\{w_{L},w_{H}\}<w_{i}$ F F Any $y_{L}<y_{i}<\min\{y_{H},w_{H}\}$ K $(1-q(w))(1-q(w_{L}))$
Notes: F = failed, S = success, K = keep. We ignore the equality case like $w_{L}=y_{i}$ in the probability because the body weight information works as a tie-breaker for the case.
Given the values of $w_{L},w_{H},y_{L},y_{H}$, and $y_{i}$ with the rank-dependent value function in Section 3, the expected evaluation of the lottery $(w_{i},q(w_{i}))$ for lifter $i$ with reference point $r_{i}$ is
$EU_{i}=v(x_{i}-r_{i})q(w_{i})$
=$=\left\{\begin{array}[]{ll}\Bigl{(}q(w_{i})q(w_{L})q(w_{H})+q(w_{i})q(w_{L})(1-q(w_{H}))P(w_{i}<y_{H})+(1-q(w_{i}))q(w_{L})\Bigr{)}[-\lambda]&\text{if }w_{i}<w_{L},w_{H}\\ +\Bigl{(}q(w_{i})(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(w_{i}))q(w_{L})q(w_{H})P(w_{L}>y_{i})+(1-q(w_{i}))q(w_{L})(1-q(w_{H}))P(w_{L}>y_{i})\Bigr{)}[-\lambda]&\text{if }w_{L}<w_{i}<w_{H}\\ +\Bigl{(}q(w_{i})q(w_{L})(1-q(w_{H}))P(w_{i}>y_{H})+q(w_{i})(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(w_{i}))q(w_{L})P(w_{L}>y_{i})\Bigr{)}[-\lambda]+q(w_{i})[1]&\text{if }w_{L},w_{H}<w_{i}\end{array}\right.$ (8)
=$=\left\{\begin{array}[]{ll}\Bigl{(}q(w_{i})q(w_{L})q(w_{H})+q(w_{i})q(w_{L})(1-q(w_{H}))P(w_{i}<y_{H})+(1-q(w_{i}))q(w_{L})\Bigr{)}[-\lambda]&\text{if }w_{i}<w_{L},w_{H}\\ +\Bigl{(}q(w_{i})(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(w_{i}))q(w_{L})P(w_{L}>y_{i})\Bigr{)}[-\lambda]+\Bigl{(}q(w_{i})(1-q(w_{H}))P(w_{i}>y_{H})\Bigr{)}[1]&\text{if }w_{L}<w_{i}<w_{H}\\ \Bigl{(}(1-q(w_{i}))q(w_{L})P(w_{L}>y_{i})\Bigr{)}[-\lambda]+q(w_{i})[1]&\text{if }w_{L},w_{H}<w_{i}\end{array}\right.$ (13)
Although the representation of the expected evaluation of the lottery in Equation (13) may appear complex, in particular, when $w_{L}<w_{i}<w_{H}$, it reveals that there are discontinuities at $w_{L}$ and $w_{H}$. Additionally, it is crucial to note that $P(w_{i}>y_{H})$ (that is, chance of the higher rival’s being turned around by the lifter) and $P(w_{L}>y_{i})$ (that is, chance of the lower rival’s turning around the lifter) are key factors that capture the pressure exerted by lower and higher rivals, respectively.
Even without further specification, Equation (13) offers theoretical predictions regarding the optimal choice of attempt weight $w_{i}$ for each domain of $w_{i}$. First, if lifter $i$ aims to choose $w_{i}$ such that $\max\{w_{L},w_{H}\}<w_{i}$, the optimal choice for $w_{i}$ is $w_{H}+\kappa$, where $\kappa$ is the minimum weight increment as specified by the competition rules, given that $q(w_{i})$ is a decreasing function of $w_{i}$. Second, if lifter $i$ chooses $w_{i}$ such that $w_{i}<\min\{w_{L},w_{H}\}$, the optimal choice is the $w_{i}$ that maximizes $q(w_{i})P(w_{i}>y_{H})$, which leads to $w_{i}=y_{H}+\kappa$. Third, if lifter $i$ selects $w_{i}$ such that $w_{L}<w_{i}<w_{H}$, the optimal choice is $w_{i}$ that maximizes $q(w_{i})(1-q(w_{H}))P(w_{i}>y_{H})$, which also results in $w_{i}=y_{H}+\kappa$. Further analysis to precisely determine the analytical optimal attempt weight $w_{i}$ across all domains of attempt weights would require additional specifications for $q(\cdot),q(w_{H}),q(w_{L}),\lambda$, and other parameters, which increases the complexity of the modeling, identification, and estimation processes, and is beyond the scope of this study and our data.